Dorcan Corporation manufactures and sells T-shirts imprinted with college names and slogans. Last year, the shirts sold for $7.50 each, and the variable cost to manufacture them was $2.25 per unit. The company needed to sell 20,000 shirts to break-even. The after tax net income last year was $5,040. Donnelly's expectations for the coming year include the following: (CMA adapted) The sales price of the T-shirts will be $10. Variable cost to manufacture will increase by one-third. Fixed costs will increase by 10%. The income tax rate of 40% will be unchanged. Based on a $10 selling price per unit and if Dorcan Corporation wishes to earn $37,800 in after tax net income for the coming year, the company's sales volume in dollars must be: Multiple Choice $213,750. $257,625. $207,000. $255,000.
Answer #1
Break-even sales 20000
X Unit Contribution margin 5.25 =7.5-2.25
Fixed costs 105000
Coming year:
Variable cost per shirt 3.00 =2.25+(2.25*1/3)
Fixed costs 115500 =105000*1.10
Required before tax income 63000 =37800/(1-40%)
Add: Fixed costs 115500
Required Contribution margin 178500
Divide by Contribution margin per shirt 7 =10-3
Required sales in units 25500
Required sales volume in dollars 255000 =25500*10
Option D $255,000 is correct